OpenStudy (inhaleof):

Which of the following is not an equivalent form of the compound inequality x + 6 < 9 or x + 6 greater than or equal to 15 x < 3 or x greater than or equal to 9 A number line with an open circle on 3, a closed circle on 9, and shading in between. x greater than or equal to 9 or x < 3 A number line with an open circle on 3, shading to the left, and a closed circle on 9, shading to the right.

OpenStudy (texaschic101):

x + 6 < 9 x + 6 >= 15 x < 9 - 6 x >= 15 - 6 x < 3 x >= 9 the incorrect one is : A number line with an open circle on 3, a closed circle on 9, and shading in between any questions ?

OpenStudy (texaschic101):

The reason that this statement is not equivalent is because it is saying that x is greater then 3 and less then 9 and that is not true. X is less then 3 and greater or equal to 9.

OpenStudy (inhaleof):

thank you very much

OpenStudy (texaschic101):

just remember this... if there is an equal sign, it is a closed circle if there is no equal sign, it is an open circle < if less then, shade to the left > if greater then, shade to the right